Question 12

  1. State the Heckscher-Ohlin Theorem.
  2. Suppose there are two countries, USA and Germany. USA is regarded as the home country and Germany is the foreign country. USA has 100 units of labor available and Germany has 80 units of labor. Both countries can produce only two goods, airplanes and cars. The output per hour of labor in the production of airplanes in the USA is 12, while in car production the output per hour of labor is 6. In Germany, the output per hour of labor in the production of airplanes is 2 and the output per hour of labor in the production of cars is 2. Using the Ricardian model,

  1. Construct a graph of the United States production possibility frontier (use the vertical axis to represent airplanes and the horizontal axis to represent cars).
  2. Construct a graph of Germany’s production possibility frontier.
  3. Suppose the mutually beneficial terms of trade between USA and Germany is 1 airplane for ¾ car, draw the consumption possibility frontier for USA by superimposing it on the graph you constructed in (i) above.
  4. Using the inverse of the marginal product of labor, construct a table of the unit labor requirement?
  5. Derive the opportunity cost for both USA and Germany for both goods using the unit labor requirement table that you obtained in (iv) above. (5 points)
  6. Based on your results, determine which country should export airplane or cars.

Question 13

  1. Why is the overall cost of living in China lower than in Europe, the United States and Canada? (7 points)

  1. Despite major gains, Chinese manufacturing workers have much lower productivity than their U.S counterparts. Chinese service workers are relatively more productive, but most services aren't tradable. So, which matters for Chinese wages-manufacturing or service productivity?
  2. Mexico and Brazil have different trading partners. While Mexico trades mainly with the United States, Brazil trades about equally with the United States and with the EU (European Union). In addition, Mexico does much more trade relative to its GDP. Explain these differences using the gravity model of trade.
  3. Explain clearly the analogy between international borrowing and lending and ordinary international trade.
Answer #1

a) Heckscher-Ohlin theorem states that a country must produce a good which requires more of its abundant resources i.e., a capital abundant country will produce and export a capital-intensive good whereas, a labor abundant country will produce and export a labor-intensive good.

b. i) USA has 100 units of labor available whereas, Germany has 80 units of labor available.

Now, output per hour of labor in USA is 12 units of airplane or 6 units of car.

Similarly, output per hour of labor in Germany is 2 units of airplane or 2 units of car.

Using the given information, we can construct production possibility frontier for USA and Germany as follows:

Here, airplanes are plotted on vertical axis and cars are plotted on the horizontal axis.

Opportunity cost of producing airplanes in USA = units of cars sacrificed/units of airplanes produced = 6/12 = 0.5 and opportunity cost of producing cars in Germany = units of airplanes sacrificed/units of cars produced = 12/6 = 2

Similarly, opportunity cost of producing airplanes in Germany = units of cars sacrificed/units of airplanes produced = 2/2 = 1 and opportunity cost of producing cars in Germany = units of airplanes sacrificed/units of cars produced = 2/2 = 2

Thus, USA has comparative advantage in the production of airplanes (lower opportunity cost of airplanes for USA) and Germany has comparative advantage in the production of cars (lower opportunity cost of cars for Germany).

ii) We have shown PPF for Germany in the above diagram.

iii)

Initially, let USA's consumption point was A i.e., 6 units of airplanes and 3 units of cars. Now, for trade, USA will produce 12 units of airplanes only (according to its comparative advantage in airplanes).

If USA consumes 6 units of airplanes, it can export another 6 units of it to Germany and in return it gains 6*(3/4) = 4.5 units of cars. Thus, the new consumption point for USA is B. We can derive the consumption possibility frontier for USA as shown in the diagram above.

Similarly, let Germany's consumption point was C i.e., 1 unit of airplane and 1 unit of car. Now, for trade, Germany will produce 2 units of cars only (according to its comparative advantage in cars).

If Germany consumes 1 unit of car, it can export another 1 unit of it to USA and in return it gains 1/(3/4) = 4/3 = 1.3 unit of airplanes. Thus, the new consumption point for Germany is D. We can derive the consumption possibility frontier for Germany as shown in the diagram above.

iv) Unit labor requirement can be constructed as follows:

v) Opportunity cost of producing airplanes for USA = unit labor requirement for airplanes / unit labor requirement for cars = (100/12)/(100/6) = 0.5

and opportunity cost of producing cars for USA = unit labor requirement for cars / unit labor requirement for airplanes = (100/6)/(100/12) = 2

Opportunity cost of producing airplanes for Germany = unit labor requirement for airplanes / unit labor requirement for cars = (80/2)/(80/2) = 1

and opportunity cost of producing cars for Germany = unit labor requirement for cars / unit labor requirement for airplanes = (80/2)/(80/2) = 1

vi) As the opportunity cost of producing airplanes is lower for USA, it has comparative advantage in the production of airplanes and thus, it will produce and export airplanes.

Again, as the opportunity cost of producing cars is lower for Germany, it has comparative advantage in the production of cars and thus, it will produce and export cars.